OpenStudy (anonymous):

(x+y)^2/2 * 60/5(x+y)

OpenStudy (anonymous):

is this \[\frac{(x+y)^2}{2}\times \frac{60}{5(x+y)}\]?

OpenStudy (anonymous):

if so do some canceling and get \[6(x+y)\]

OpenStudy (anonymous):

yes thats the equation but it should end in a fraction